# Image Embedding Sampling Method for Diverse Captioning

### v1

@@ 4 Results
HBoP achieves the best diversity scores while maintaining relevance among smaller VLMs. We compare the diversity and relevance of captions generated by different models, as shown in Table 1. While larger VLMs with LLM-based text encoders generally achieve stronger overall performance, HBoP remains competitive, achieving the closest diversity scores to the gold captions among smaller VLMs in terms of PCD (see Appendix C for details), mBLEU-4 and Div-2 Aneja et al. (2019b).
Specifically, compared to captions generated by BLIP Li et al. (2022a) and BLIP-2 Li et al. (2023), HBoP achieves a PCD score closest to that of the gold captions, along with the lowest mBLEU-4 and highest Div-2 scores among BLIP (both with and without nucleus sampling (NS) Li et al. (2022a)), Seq-CVAE Aneja et al. (2019b), and ModeCAP [^1] The dataset annotations and features necessary to train ModeCap are exclusively available for the MSCOCO dataset, making it difficult to replicate the experiments for fair comparison on the NoCaps and Flickr30k datasets. Chen et al. (2022). Notably, in some cases, HBoP also surpasses larger VLMs with LLM-based encoders, such as BLIP-2 Li et al. (2023), Honeybee-7B Cha et al. (2023), and LLaVA-1.5 Liu et al. (2023a), in diversity metrics, as indicated by lower mBLEU-4 and higher Div-2 scores (notice red mBLEU-4 and blue Div-2, marking stronger diversity performance for HBoP).
At the same time, HBoP maintains strong similarity between generated captions and reference texts, as measured by SBERT Reimers and Gurevych (2019), and preserves high image-text alignment, as indicated by CLIP-Score Hessel et al. (2021). Notably, HBoP achieves SBERT and CLIP-Score values that are comparable to BLIP, BLIP-NS, and LLaVA, while outperforming HoneyBee in both metrics. Although BLIP-2 scores the highest, HBoP remains competitive, demonstrating a strong balance between relevance and diversity.
HBoP generates semantically meaningful captions. We evaluate the semantic integrity of HBoP captions against those generated by other models, using LLama-2-13b Touvron et al. (2023) and GPT-4 Fu et al. (2023). As shown in Table 2, HBoP achieves semantic integrity scores close to the gold captions. Notably, HBoP outperforms models like PnP-VQA in this metric. We attribute this surpass to our method, which samples more meaningful image embeddings using the proposed HCM component.
5 Conclusion
6 Limitations
The HBoP architecture uses bounding box information to extract image embeddings, where the bounding boxes may contain additional information beyond the segmented object. The exploration of caption generation solely based on irregular-shaped segmentation masks is deferred as future work.
7 Ethical Statement
Captions generated with HBoP might inadvertently contain harmful content. However, the final caption outputs mainly depend on the image content and pretrained image captioning model. Therefore, unless the images themselves are harmful or the pretrained model produces unsafe captions, HBoP captions are expected to pose minimal risk.
Appendix A Appendix
Appendix B Additional Related Works
B.1 Vision-Language Models (VLMs)
A growth of interest in VLMs has continued due to the wide availability of multi-modal data on the web. Foundation VLMs can be applied to a range of tasks in a zero-shot manner. Notably, CLIP Radford et al. (2021b) jointly pre-trains an image encoder and a text encoder by maximizing and minimizing the cosine similarity of correct and incorrect image-text pair embeddings respectively with image-text contrastive (ITC) loss. In contrast, BLIP Li et al. (2022a) uses both ITC and image-text matching (ITM) loss for enhanced image-text data representation. Additionally, the BLIP Li et al. (2022a) captioner uses language modeling (LM) loss for autoregressive image caption generation along with a filter, capfilt to improve the quality of image-text pairs for training.
Flamingo Alayrac et al. (2022) shows remarkable zero-shot ability in image captioning, visual question-answering (VQA), and image-text retrieval (ITR) tasks by leveraging the few-shot learning ability of pre-trained vision-only and language-only models. It simply interleaves input visual data with task-specific text examples, producing free-form texts for unseen visual data. Another general-purpose model, BEIT3 Wang et al. (2022b) with Multiway Transformer structure, uses different types of modality experts to perform fusion and modality-specific training. A masked modeling objective on images only and image-text pairs is performed for computer vision tasks (e.g., image classification, semantic segmentation, object detection) and vision-language tasks (e.g., VQA), respectively. Whereas the VQA task uses a fused encoder for image-text pairs, the ITR task encodes images and texts independently with ITC loss. Lastly, sequence-to-sequence learning is applied to generate texts from images for the image captioning task. Inspired by these previous works, we propose a meta-VLM model that utilizes a pre-trained BLIP Li et al. (2022a) image captioning module to generate enhanced textual representations, which can later serve as useful data for various downstream tasks.
B.2 Hierarchical Representation
Identifying and extracting regions of interest within images is crucial for a hierarchical representation. The most intuitive way to achieve this would typically involve the use of object detectors Yao et al. (2019); Cornia et al. (2020); Zhang et al. (2021). However, the heavy computational demands of the object detectors inevitably lead to inefficiency during the inference stage Yao et al. (2019); Cornia et al. (2020); Zhang et al. (2021). In response, recent works sought to replace these cumbersome detectors by adopting visual concepts in the form of object tags Fang et al. (2022); Shukor et al. (2022) as an alternative. However, this detector-free approach is contingent upon the availability of object-specific data within the dataset. Employing pre-trained models is a more efficient way to identify areas of interest within images. GradCAM Selvaraju et al. (2017) highlights essential regions that the pre-trained models used to predict any target concept using its gradients with respect to feature map activations of the final convolutional layer. DINOv2 Oquab et al. (2023) capitalizes on existing self-supervised pre-trained models to generate robust, all-purpose visual features, supporting a wide array of tasks ranging from image-level classification to pixel-level segmentation. However, the image regions/features delineated by GradCAM/DINOv2 tend to show saliency for specific tasks and are unable to capture the full spectrum of visual representations. Conversely, SAM Kirillov et al. (2023) intricately segments every semantically significant component of an image into high-quality segmentation masks generated by prompting with various inputs such as point, box, mask, or free-form text, unrestricted with types of tasks. In our framework, we integrate SAM Kirillov et al. (2023) to create semantically meaningful segmentation masks for an entire image automatically.
Several prior studies have incorporated the principles of hierarchy or multi-scale representation into their model architectures, aiming to enhance the alignment between images and texts Ji et al. (2021); Shao et al. (2023); Shukor et al. (2022). SHAN Ji et al. (2021) deconstructs the image-text matching process into two distinct facets: fragment-level and context-level alignments enabling matches across three different scopes: local-to-local, global-to-local, and global-to-global. HiVLP Shao et al. (2023) leverages both low- and high-dimensional features to represent coarse and fine details. ViCHA Shukor et al. (2022) aligns images and texts across various layers of neural network encoders with the underlying assumption that each layer reflects varying semantic levels. Unlike these approaches, we divide the segmentation masks hierarchically and use the embeddings of the extracted individual image patches for caption generation.
B.3 Caption Evaluation
Common image captioning evaluation metrics, including BLEU Papineni et al. (2002a), METEOR Banerjee and Lavie (2005), ROUGE Lin (2004), and CIDEr Vedantam et al. (2015) scores are primarily n-gram approaches that assess the quality of generated captions by considering their overlap with human-generated captions. Most SOTA VLMs frequently exhibit promising scores across these conventional evaluation metrics. However, these metrics are limited in their capabilities to measure the diversity of the generated captions. This limitation leads to a bias in these models towards generating an "average" and "safe" caption reflecting the most basic information in the image, rendering them less informative than human-generated captions. To address this gap, we incorporate several diversity metrics, including mBLEU-4, Div-2 Aneja et al. (2019b), and the proposed pairwise cosine distance (PCD), along with semantic integrity and relevance scores to ensure that the captions generated by our framework are not only diverse but also meaningful and directly relevant to the given image and human-annotated captions.
Appendix C Experiments
C.1 Implementation Details
The ISM (Section 3.1) employs the fully automated SAM with no prompting Kirillov et al. (2023), along with the image encoder initialized from ViT (ViT-L/16) pre-trained on ImageNet Dosovitskiy et al. (2021), following the same settings as BLIP Li et al. (2022a). Note that we use BLIP Li et al. (2022a) for captioning instead of BLIP-2 Li et al. (2023) since BLIP-2 uses intermediate representations trained on pairs of entire images and texts for caption generation using an LLM, which is not directly applicable to HBoP that uses pairs of image patches and texts. The HCM (Section 3.2) creates the global level by selecting the top ($k=$) 5 masks with the largest areas and designating the remaining masks as fine-grained. To create the regional level, K-means clustering, with ($K=$) 5 clusters per image, is applied to the bounding boxes of the segmentation masks. NMS with a threshold of 0.1 is applied at all three levels. Lastly, the ICM (Section 3.3) follows the methodology outlined in Tiong et al., 2022.
All the model captions in Tables 1 and 3 are regenerated, except for Seq-CVAE Aneja et al. (2019b), where the results are taken directly from the original paper. While HBoP benefits from bounding box information, it is important to note that other baseline methods (e.g., ModeCap) have the additional advantage of explicit learning objectives to improve diversity. The exact prompts we use for Honeybee Cha et al. (2023) (top) and LLaVA-1.5/1.6 Liu et al. (2023a) are in Table 4.
C.2 Evaluation
C.2.1 Diversity
We measure the diversity in the generated captions using the cosine similarity between the sentence embeddings of all the corresponding captions per image. The comparison baselines are random captions, where each caption corresponds to different images, BLIP Li et al. (2022a) with and without nucleus sampling (NS [^2] Unless otherwise specified, all the BLIP models in this paper refer to BLIP with NS.) Holtzman et al. (2019), BLIP-2 Li et al. (2023), ModeCap Chen et al. (2022), Honeybee Cha et al. (2023), and gold captions [^3] We exclude PnP-VQA since the captions are generated per question instead of per image, unlike other baselines.. The diversity of the generated captions ($s_{1},s_{2},...s_{n}$) per dataset instance [^4] Note that $n=5$ for all dataset instances, and we use one global caption and five regional captions for HBoP. is measured using pairwise cosine distance (PCD):
In the above equation, $cos$ represents the cosine similarity of the input embeddings. We use sentence embeddings from a pre-trained sentence transformer model (all-MiniLM-L6-v2) Reimers and Gurevych (2019), denoted as $M$ in the Eq. 1 that can capture the semantic relationships between captions. This measure evaluates the extent to which the generated captions differ from each other per image. We report the final diversity score for each dataset as the averaged PCD scores of all images in the dataset. Ideally, the PCD score should be lower than that of random captions that serve as the upper bound of the diversity score, but it should be higher than that for captions generated by existing baselines.
C.2.2 Relevancy
While confirming that each dataset contains captions with high semantic integrity is crucial, the captions must also be relevant to the corresponding images. We employ CLIP-Score Hessel et al. (2021) that calculates the correlation between visual and textual CLIP embeddings Radford et al. (2021b) using pre-trained ViT (openai/ clip-vit-base-patch32) without relying on human-generated references. Similar to the comparison baseline datasets for semantic integrity evaluation, we compare HBoP with PnP-VQA Tiong et al. (2022), BLIP Li et al. (2022a), BLIP-2 Li et al. (2023), gold captions, along with random captions. We generate random captions by selecting five random captions for each image from a pool of HBoP captions corresponding to different images. In other words, although the random caption itself should make sense, they depict mismatched images. We randomly select one out of a total of five captions per image for each dataset and compute the correlation between CLIPScores of generated captions and gold captions.
C.2.3 Semantic Integrity
We measure the semantic integrity of the generated captions (Table 2) to ensure meaningfulness and cohesiveness using LLM evaluation as it has empirically shown high correlation with human judgment Chiang and yi Lee (2023); Liu et al. (2023b); Fu et al. (2023). We prompt Llama-2-13B (Llama-2-13b-chat -hf) Touvron et al. (2023) to access the semantic integrity of HBoP captions along with gold and other baselines (PnP-VQA Tiong et al. (2022), BLIP Li et al. (2022a), BLIP-2 Li et al. (2023)) captions. Specifically, we randomly select two captions out of a total of five captions per image for each dataset and evaluate the semantic integrity by averaging the coherency and meaningfulness scores for each caption using the prompt shown in Table 5. We use the prompt "This is a picture of" to generate captions for all models in our experiments. This deliberate choice ensures a fair comparison of the general caption generation ability across models, as altering the prompt can yield significantly different results, making fair evaluation challenging.
@@ D.1 Relevancy
In Figure 3, HBoP captions (y-axis values in the last column) show comparable relevance scores with gold captions (x-axis values in the last column) with the slope of a linear regression line [^5] The p-values for all the regression lines are less than 0.001, except for the those of lines in the first columns, which are not statistically significant being close to 0.5. Although the slopes of these regression lines (MSCOCO Lin et al. (2014): 0.42, Flickr30k Young et al. (2014): 0.39, Nocaps Agrawal et al. (2019): 0.34) are less than those of BLIP Li et al. (2022a) (0.49, 0.44, and 0.45) and BLIP-2 Li et al. (2023) (0.51, 0.45, 0.43), we observe a trend of having relevance scores in the range of 20 to 40 for both x and y axes values. On the other hand, relevance scores for random and PnP-VQA Tiong et al. (2022) captions have a spurious and less-correlated relation with those of gold captions.
@@ D.3 Fine-grained Captions
Although not evaluated in the perspectives of three main evaluation metrics, we can also create what we refer to as fine-grained captions that can serve as image tags using our proposed methodology. These serve as supplementary information, enhancing the depth of understanding of the image. They are more vital when dealing with complex images containing various small or intricate objects, which conventional caption generation processes may often overlook. By introducing the additional layer of granularity, our approach ensures a more detailed and inclusive interpretation of the image.
@@
[INST] <<much-less-than$<$ <SYS>>much-greater-than$>$ > You will be given a caption generated from an image. Given criteria and rating options, rate the response. Respond with a number only. Evaluation Criteria: [CRITERION]: [DEFINITION] Scale: from 1 to 10 Answer: <<much-less-than$<$ </SYS>>much-greater-than$>$ > INPUT [///INST] [CRITERION]: Coherence/Meaningfulness [DEFINITION]: the logical and clear connection between ideas or elements within a context. It is characterized by the consistency, integrity, and clarity of information or arguments presented./the relevance and significance of the content in the caption. A meaningful caption goes beyond a literal description, providing insight, context, or emotion that enhances the viewer’s understanding or appreciation of the image.

### v2

@@ D.1 Relevancy
In Figure 3, HBoP captions (y-axis values in the last column) show comparable relevance scores with gold captions (x-axis values in the last column) with the slope of a linear regression line‡‡‡‡‡‡The p-values for all the regression lines are less than 0.001, except for the those of lines in the first columns, which are not statistically significant being close to 0.5. Although the slopes of these regression lines (MSCOCO Lin et al. (2014): 0.42, Flickr30k Young et al. (2014): 0.39, Nocaps Agrawal et al. (2019): 0.34) are less than those of BLIP Li et al. (2022a) (0.49, 0.44, and 0.45) and BLIP-2 Li et al. (2023) (0.51, 0.45, 0.43), we observe a trend of having relevance scores in the range of 20 to 40 for both x and y axes values. On the other hand, relevance scores for random and PnP-VQA Tiong et al. (2022) captions have a spurious and less-correlated relation with those of gold captions.
@@ D.3 Fine-grained Captions
Although not evaluated in the perspectives of the three main evaluation metrics, we can also create what we refer to as fine-grained captions that can serve as image tags using our proposed methodology. These serve as supplementary information, enhancing the depth of understanding of the image. They are more vital when dealing with complex images containing various small or intricate objects, which conventional caption generation processes may often overlook. By introducing the additional layer of granularity, our approach ensures a more detailed and inclusive interpretation of the image.
D.4 FastSAM Results
While the segmentation step introduces additional overhead, the hierarchical embedding process itself does not, as it operates directly on the single full-image embedding. Therefore, embedding is not a computational bottleneck in our pipeline. As shown below (Table 5), we provide a comparison of inference time (average per image) between BLIP and HBoP on the MSCOCO dataset. While SAM does increase inference time, it brings significant gains in diversity without sacrificing performance. Furthermore, this overhead can be drastically reduced by replacing SAM with FastSAM Zhao et al. (2023), which provides comparable segmentation quality at 50–170× faster runtime.
Models Segmentation Time (s/img) Captioning Time (s/caption)
BLIP 0.38
HBoP (SAM) 5.43 0.38
HBoP (FastSAM) 0.18 0.38
Table 5: Average segmentation and captioning time for different models.
@@ D.4 FastSAM Results
[INST] <<<<SYS>>>> You will be given a caption generated from an image. Given the criteria and rating options, rate the response. Respond with a number only. Evaluation Criteria: [CRITERION]: [DEFINITION] Scale: from 1 to 10 Answer: <<<</SYS>>>> INPUT [//INST] [CRITERION]: Coherence/Meaningfulness [DEFINITION]: the logical and clear connection between ideas or elements within a context. It is characterized by the consistency, integrity, and clarity of information or arguments presented./the relevance and significance of the content in the caption. A meaningful caption goes beyond a literal description, providing insight, context, or emotion that enhances the viewer’s understanding or appreciation of the image.